formic acid, HCOOC, can decompose to fom carbon monoxide gas, CO, and liquid water. if 3.85 L of CO was collected by downward displacement of water at 25.0 degree celcius and 689 mmHg how many grams of formic acid were reacted


1
Expert's answer
2021-06-01T02:44:55-0400

Ans:- Reaction of Decomposition of Formic Acid "(HCOOH)"


"HCOOH \ \ \ \to \ \ \ CO(gas)\ \ \ + (liq. )H_2O"


Carbon Mono Oxide was collected at temperature "(T)""=25+273=298K" and at Pressure"(P)=689\ mmHg" and having Volume"(V)=3.85 \ L"

Value of Gas Constant"\ =62.363\ mmHg.L.K^{-1}⋅mol^{−1}"


Ideal Gas equation

"PV=nRT\\"

"n=\dfrac{PV}{RT}"



"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ n=\dfrac{689\times3.85}{62.363\times298}=0.1427\ mol"


"0.1427 \ mol" of Carbon Mono Oxide was formed by the decomposition of formic acid "(HCOOH)"

So, According to Stoichiometry

"1\ mol" of formic acid react to form = "1\ mol" of Carbon Mono Oxide

Hence "0.1427 \ mol" of formic Acid "(HCOOH)" was reacted

Therefore Gram of Formic Acid "(HCOOH)" react"=0.1427\times46=6.5642\ gm"